OpenStudy (anonymous):

2. Simplify the expression –4(x – 7) + 3(x + 9). Need helpe can someone help me

OpenStudy (zzr0ck3r):

first distribute the -4 and the 3 \(a(b+c)=ab+ac\)

OpenStudy (zzr0ck3r):

then tell me what you got

OpenStudy (anonymous):

7

OpenStudy (zzr0ck3r):

? show me how you got 7

OpenStudy (anonymous):

4+3=7

OpenStudy (zzr0ck3r):

I said to distribute, not add them together... \(-4(x-7)+3(x+9)=-4x+28+3(x+9)\) I distributed the -4, now do something similar with \(3(x+9)\) and show me what you got.

OpenStudy (anonymous):

i am confused i am not good with this stuff

OpenStudy (zzr0ck3r):

if I tell you that \(a(b+c)=ab+ac\) then \(3(x+9)=?\)

OpenStudy (anonymous):

3x+9?

OpenStudy (zzr0ck3r):

3x+3*9

OpenStudy (zzr0ck3r):

do you see how the 3 gets multiplied by the x and the 9?

OpenStudy (zzr0ck3r):

this is very important.

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

ok, so we have \(-4(x-7)+3(x+9)=-4x+28+3(x+9)=\\ -4x+28+3x+27\)

OpenStudy (zzr0ck3r):

you with me

OpenStudy (zzr0ck3r):

?

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

now group the like terms \(-4x+3x+28+27\)

OpenStudy (zzr0ck3r):

can you simplify that?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

what do you got?

OpenStudy (anonymous):

54

OpenStudy (zzr0ck3r):

what is \(-4x+3x\) ?

OpenStudy (anonymous):

-1

OpenStudy (zzr0ck3r):

the x does not go away

OpenStudy (anonymous):

well then i dont know

OpenStudy (zzr0ck3r):

Ok what about 3x+4x?

OpenStudy (zzr0ck3r):

if you have 3 of something and you add 4 more, then you have 7 of that something the something here is x, so 3x+4x=7x make sense?

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

So \(-4x+3x = -x\)

OpenStudy (zzr0ck3r):

or \(-1*x\) (same thing(

OpenStudy (anonymous):

but i need a answer

OpenStudy (zzr0ck3r):

But we just write it as \(-x\)

OpenStudy (zzr0ck3r):

ok so \(-4x+3x + 28+27 = -x+55\)

OpenStudy (anonymous):

ok